ITF found in LOW_NOISE_3_SQZ. Unlocked at 16:41 UTC, it wasn't initially possible to lock the OMC due to the Peltier disabled. I contacted the DET Oncall ( M. Eisenmann ) that started investigating the issue at LOW_NOISE_1, but at 17:37 UTC the ITF unlocked again and we started experiencing some strange behavior of the B7 signals ( see entry logbook #45474 ). This issue was solved at 19:33 UTC and it was possible to relock again, ITF back in Science Mode at 19:59 UTC.
ITF left locked.
ITF_LOCK, OMC_LOCK and DET_MAIN crashed after the unlock at 16:41 UTC. Properly restarted.
ITF temporary exited LOW_NOISE_3_SQZ at 15:21 UTC, the Automation set it back to the same state a few seconds later. Back in Science mode at 15:22 UTC.
Downtime Event - Detection
ITF in TROUBLESHOOTING Mode from 17:07 UTC to 19:59 UTC. Initially, the mode was set due to the OMC peltier disabled after the unlock at 16:41 UTC, the period was extended because of the issue with the B7 signals.
14:43 UTC - 15:19 UTC
16:49 UTC - 17:27 UTC
18:57 UTC - 19:21 UTC
21:10 UTC - 21:40 UTC